Data for Barry Computer Co. and its industry averages follow. The firms debt is priced at par, so the market value of its deBarry Computer Company: Income Statement for Year Ended December 31, 2019 (In Thousands) $2,150,000 Sales Cost of goods solda. Calculate the indicated ratios for Barry. Do not round intermediate calculations. Round your answers to two decimal placesPlease and thank you
Answer #1

a

Current ratio = current assets/current liablities
Current ratio = 875265/418605
Current ratio = 2.09
Quick or acid test ratio = (current assets-inventory)/current liablities
Quick or acid test ratio = (875265-253700)/418605
Quick or acid test ratio = 1.48
Receivables turnover = Credit sales/receivables
Receivables turnover = 2150000/431290
Receivables turnover = 4.99
days of sales outstanding = number of days in a year/receivables turnover
days of sales outstanding = 365/4.99
days of sales outstanding = 73.15
Inventory turnover = COGS/inventory
Inventory turnover = 1913500/253700
Inventory turnover = 7.54
Please ask remaining parts seperately,